eriodic Inventory Using FIFO, LIFO, and Weighted Average Cost Methods The units of an item available for sale during the year were as follows: Jan. 1 Inventory 4 units at $50 $200 Aug. 7 Purchase 17 units at $51 867 Dec. 11 Purchase 12 units at $52 624 33 units $1,691 There are 19 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost using (a) the first-in, first-out (FIFO) method; (b) the last-in, first-out (LIFO) method; and (c) the weighted average cost method (round per unit cost to two decimal places and your final answer to the nearest whole dollar). a. First-in, first-out (FIFO) $ 988 b. Last-in, first-out (LIFO) $ 969 c. Weighted average cost $ 950 Feedback
Answer #1
Unit Unit price Total Cost
Jan 1 4 50 200
Aug 7 17 51 867
Dec 11 12 52 624
Total 33 1691

Calculate ending inventory

FIFO LIFO Weighted average
Ending inventory (12*52+7*51) = 981 (4*50+15*51) = 965 1691/33*19 = 974